# Error Estimate and Convergence Analysis for Data Valuation

### v2

@@ 1 Introduction
  • Provides the first rigorous error and convergence guarantees for the NDDV method by formulating data valuation as a discrete-time stochastic optimal control problem and exploiting its Hamiltonian structure for analysis.
  • Establishes quantitative stability/error estimates for loss differences under standard Lipschitz and smoothness assumptions, yielding quadratic bounds with respect to control perturbations and clarifying the effect of time discretization on accuracy.
  • Proves that the expected squared gradient norm of the training loss vanishes asymptotically under appropriate step-size conditions, ensuring asymptotic stationarity of the NDDV training dynamics.
  • Demonstrates that the associated meta-level optimization for an $L$-smooth meta loss attains a sublinear convergence guarantee of order $O(1/\sqrt{K})$, implying $\varepsilon$-stationarity in $O(1/\varepsilon^{2})$ iterations.
@@ 2.5 Regularity Assumptions
$\|\mathcal{V}(\cdot)\bigl(\Phi(X)-\Phi(X^{\prime})\bigr)\|+\|\mathcal{V}(\cdot)\bigl(\nabla\Phi(X)-\nabla\Phi(X^{\prime})\bigr)\|\leq K\|X-X^{\prime}\|$
@@ 3.3 Main Stability Estimate
$\|\Delta X_{i,t}\|+\|\Delta Y_{i,t}\|\leq C\sum_{s=0}^{T-1}\|\psi_{i,s}-\psi^{\prime}_{i,s}\|.$
@@ 4.3 Convergence of the Meta Loss
$\qquad+\tfrac{L\sigma^{2}}{2}(\beta^{k})^{2}.$
@@
$\sum_{k=1}^{K}\Bigl(\beta^{k}-\tfrac{L}{2}(\beta^{k})^{2}\Bigr)\mathbb{E}\|\nabla\ell(\hat{\psi}^{k}(\theta^{k}))\|^{2}\leq C$
@@
Figure 1: Convergence behavior of NDDV. a. Training loss curves. b. Meta loss curves. Both curves illustrate the monotone decay and sublinear convergence predicted by the theory.
@@ 5 Convergence Verification
Figure 2: State behavior of NDDV. a. State trajectories. b. Co-state trajectories. Both trajectories illustrate the evolution of each data point’s state under the SMP.
@@ 5.2 Results
Figure 2(a) shows that the data-state trajectories evolve from being initially concentrated to becoming increasingly dispersed across all datasets. This observation suggests that NDDV progressively encodes individual data-point characteristics via the learned data-state trajectories. Similarly, the data co-state trajectories follow a consistent evolution pattern: they start highly clustered and gradually spread out as training proceeds (Fig. 2(b)). Overall, these results indicate that the data states and co-states evolve jointly, yielding trajectories that are compatible with training convergence. Moreover, according to Eq. 5, the learnable trajectories enable us to estimate the values of all data points within a single training process.

### v3

@@ 1 Introduction
  • First rigorous error and convergence analysis: We present the first formal error and convergence guarantees for the NDDV method, modeling data valuation as a discrete-time stochastic optimal control problem. By leveraging the Hamiltonian structure of the problem, we provide a comprehensive analysis.
  • Quantitative stability and error bounds: We establish precise stability and error estimates for loss differences under standard Lipschitz and smoothness conditions. Our results include quadratic bounds with respect to control perturbations and provide insights into the impact of time discretization on accuracy.
  • Asymptotic stationarity of training dynamics: We prove that, under appropriate step-size conditions, the expected squared gradient norm of the training loss converges to zero asymptotically. This ensures the asymptotic stationarity of the NDDV training dynamics.
  • Sublinear convergence guarantee: We demonstrate that the meta-level optimization associated with an $L$-smooth meta-loss achieves a sublinear convergence rate of $O(1/\sqrt{K})$, leading to $\varepsilon$-stationarity in $O(1/\varepsilon^{2})$ iterations.
@@ 2.4 Dynamic Valuation Definition
Eq 5 provides a discrete-time counterpart of the underlying continuous dynamics, yielding a unified framework that enables error estimation and convergence analysis for data valuation within a single training process. Using the Half Moons dataset as an example, NDDV is effective for standard data-valuation tasks—detecting corrupted data, removing high- and low-value data, and adding high- and low-value data (see Figure 1)—while substantially improving computational efficiency.
Fig. 1: The validity of NDDV for data valuation. On the Half Moons dataset, NDDV effectively evaluates data value in experiments on corrupted-data detection and data removal/addition.
@@ 2.5 Regularity Assumptions
$\|\mathcal{V}(\cdot)\bigl(\Phi(X)-\Phi(X^{\prime})\bigr)\|+\|\mathcal{V}(\cdot)\bigl(\nabla\Phi(X)-\nabla\Phi(X^{\prime})\bigr)\|\leq K\|X-X^{\prime}\|.$
@@ 3.3 Main Stability Estimate
$\|\Delta X_{i,t}\|+\|\Delta Y_{i,t}\|\leq C\sum_{s=0}^{T-1}\|\psi_{i,s}-\psi^{\prime}_{i,s}\|,$
@@ 4.3 Convergence of the Meta Loss
$\qquad+\tfrac{L\sigma^{2}}{2}(\beta^{k})^{2},$
@@
$\sum_{k=1}^{K}\Bigl(\beta^{k}-\tfrac{L}{2}(\beta^{k})^{2}\Bigr)\mathbb{E}\|\nabla\ell(\hat{\psi}^{k}(\theta^{k}))\|^{2}\leq C.$
@@
Fig. 2: Convergence behavior of NDDV. a. Training loss curves. b. Meta loss curves. Both curves illustrate the monotone decay and sublinear convergence predicted by the theory.
@@ 5 Convergence Verification
Fig. 3: State behavior of NDDV. a. State trajectories. b. Co-state trajectories. c. Value trajectories. Both trajectories illustrate the evolution of each data point’s state and value under the SMP.
@@ 5.2 Results
Figure 3(a) shows that the data-state trajectories evolve from being initially concentrated to becoming increasingly dispersed across all datasets. This observation suggests that NDDV progressively encodes individual data-point characteristics via the learned data-state trajectories. Similarly, the data co-state trajectories follow a consistent evolution pattern: they start highly clustered and gradually spread out as training proceeds (Figure 3(b)). Overall, these results indicate that the data states and co-states evolve jointly, yielding trajectories that are compatible with training convergence. Further, according to Eq. 5, the learnable trajectories enable us to estimate the values of all data points within a single training process. As shown in Figure 3(c), data value trajectories maintain the trends observed in both data state and co-state trajectories. These trajectories, initially nearly converging at a single point, gradually disperse over time. This result indicates that initial data points display nearly identical values. However, through the assessment by NDDV, the learned data value trajectories progressively diverge, effectively leading to a distinct valuation of data points.